OpenStudy (anonymous):

What is the graph of 4x^2 + y^2 = 16? What are itss lines o symmetry? What are the domain and range?

OpenStudy (anonymous):

\[divide~by~16\] \[\frac{ x^2 }{4 }+\frac{ y^2 }{16 }=1\] can you solve further?